Child pages
  • Commander Report Day 05 TasMars2013
Skip to end of metadata
Go to start of metadata

Date Time: 19:00 01/28/2013
Written By: David Willson, Crew Commander of MDRS123

Today was a quiet day for crew MDRS123, recovering from the long and physically active filming with Chanel Seven Australia.

The temperature was 32 degrees F in the morning rising to 50 degrees F in the afternoon. The ground was not sticky and for the first time it did not snow during the night. The weather is clearly improving compared to the previous days.

In the morning the crew undertook an out of sim EVA to survey locations for our project work. This was followed by:

- Emma checked setting on the WiFi for IT maintenance person Chuck Killan.

- David tested the Ramon and Ocean Optics spectrometers for future work and reviewed the rover control problems with rover manager Scott Davis. The rover has serious control issues to be resolved.

- Melanie continued work in the Green Hab.

In the afternoon the crew visited Gobblin Valley.

Finally we collected a new laptop for the Ramon and Ocean Optics spectrometers and outreach materials shipped from JPL.

ENDS.

  • No labels